If the first and third of 3 consecutive odd integers are added, the result is 57 less than 5 times the second integer. Find the 3rd integer.
Answer by Maths68(1474)

Let
1st Interger = x+1
2nd Integer = x+3
3rd Interger = x+5
------------------------------------------------
(x+1)+(x+5)=5(x+3)-57
x+1+x+5=5x+15-57
2x+6=5x-42
2x-5x=-42-6
-3x=-48
Divide by – 3
-3x/-3=-48/-3
-x=-16
x=16
---------------------------------------------
1st Interger = x+1=16+1=17
2nd Interger = x+3=16+3=19
3rd Interger = x+5=16+5=21
-----------------------------------------
Check
17+21=5*(19)-57
38=95-57
38=38